questions, each has 10 points. An unknown concentration of HCl is titrated using known KOH.
Answer #1

Since HCl is a strong acid and KOH is a strong base, hence it will dissociate completely into solution

The balanced reaction will be

HCl + KOH ------------- KCl + H2O

Number of moles of HCl = Number of moles of KOH

Molarity of HCl * Volume of HCl (in L) = Molarity of KOH * Volume of KOH (in L)

Molarity of HCl * 25.0/1000 = 0.125 * 43.6/1000

Molarity of HCl = (0.125 * 43.6)/(25.0) = 0.218 M

Note - Post any doubts/queries in comments section.